OpenStudy (anonymous):

if y varies jointly as x and z and y=18 when x=2 and z=3 ,, find y when x=5 and z=6

OpenStudy (anonymous):

please help

OpenStudy (mertsj):

Varies jointly means it varies directly with two or more variables. So write y = kxz.

OpenStudy (mertsj):

Then substitute the given values for y, x, and z and find k

OpenStudy (mertsj):

When you wrote x = 2 and x = 3 did you really mean x = 2 and y = 3?

OpenStudy (anonymous):

no x=2 and z=3

OpenStudy (mertsj):

18=k(2)(3)

OpenStudy (mertsj):

So what is k?

OpenStudy (anonymous):

would it be 36??? y=k/x 18=k/2 18=36/2 ?????

OpenStudy (mertsj):

Did you see the equation I wrote: y = kxz?

OpenStudy (mertsj):

That is the joint variation. y varies jointly with x and z

OpenStudy (anonymous):

im still not understandingg

OpenStudy (mertsj):

There are 3 basic types of variation. (1) y varies directly as x and the equation is y = kx (2) y varies inversely (or indirectly) as x and the equation is y = k/x (3) y varies jointly as x and z and the equation is y = kxz Which one of those is this problem you posted?

OpenStudy (mertsj):

Is it direct, inverse, or joint?

OpenStudy (anonymous):

i posted jointly but when i do jointly i have to find k and the way im doing it is the way i showed uu and i dont think when im doing jointly im doing it right

OpenStudy (mertsj):

Everything you just said is correct. So since it is joint variation, why don't you use the joint variation equation?

OpenStudy (mertsj):